OpenStudy (anonymous):

> Double checking my answer < Find the value of r in (4,r), (r,2) so that the slope of the line containing them is -5/3. So, the answer is r= -13.

jimthompson5910 (jim_thompson5910):

I'm getting a different answer. Can you post your work so I can see what went wrong?

OpenStudy (anonymous):

Give me about 3 minutes and I can.

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

Okay, so I got to -5r+20=6-3r Then I subtracted 20 from 20 and 6 so I got -5r=-26-3r. I added -5r to -5 and -3. Then I got -26 = 2r. I subtracted by 2 from each side and got r = -13.

jimthompson5910 (jim_thompson5910):

I see where you made the error. 6 - 20 is NOT equal to -26 6 - 20 is -14

OpenStudy (anonymous):

Brain fart.

jimthompson5910 (jim_thompson5910):

so you should have -5r = -14 - 3r then you add 3r to both sides to get -5r = -14 - 3r -5r + 3r = -14 - 3r + 3r -2r = -14